Two cards are drawn from a standard deck of cards. What is the probability that both cards are diamonds.
I got 1/13. Is that correct?

Answer by Alan3354(69443)

Two cards are drawn from a standard deck of cards. What is the probability that both cards are diamonds.
I got 1/13. Is that correct?
----------------------------
The 1st card is 13/52 = 1/4
If that card is not replaced,
the 2nd card is 12/51 = 4/17
(1/4)*(4/17) = 1/17